OpenStudy (anonymous):

find the distance between the two point. Round to the nearest tenth. (2,5) and (-1,-5)

OpenStudy (callisto):

use the distance formula \[\sqrt{[(2-(-1)]^2 + [(5-(-5)]^2}\]\[=\sqrt{3^2 + 10^2}\] \[=\sqrt{109}\] = 10.4 units